OpenStudy (anonymous):

The formula of the conversion degree celsius to degree farenheit is F=9/5c+32.solve for C.use this formula to find the tempreture in celsius if is 100degree F

OpenStudy (anonymous):

I want to make sure I am understanding your problem first.... F(100)=9/5c (celcius)+32?

OpenStudy (radar):

F=9/5C+32, 9/5C=F-32, C=5/9(F-32) Is this what you meant?

OpenStudy (radar):

If you understand the relationship between C and F, you can calculate a temperature where the value in C and F is the same. Hint it is cold!!

OpenStudy (radar):

O.K. C=5/9(F=32) or C=5/9(100-32) or C=5/9 * 68 =37.78